P(-2.81>z>1.05)=


1
Expert's answer
2021-10-21T12:02:52-0400

"P(-2.81<z<1.05)=P(z<1.05)-P(z<-2.81)"


From Standard Normal Distribution Table,

P(z < 1.05) = 0.85314

P(z < -2.81) = 0.00248


"\implies P(z<1.05)-P(z<-2.81)= 0.85314-0.00248 = 0.85066"

Hence, "\boxed{P(-2.81>z>1.05)=0.85066}"